# UAV-ON: A Benchmark for Open-World Object Goal Navigation with Aerial Agents

### v1

@@ 1 Introduction
Recent progress in aerial Vision-and-Language Navigation (VLN) Liu et al. (2023b); Chu et al. (2024b); Wang et al. (2024b); Gao et al. (2025); Zhao et al. (2025b) has been largely driven by detailed, fine-grained language instructions. However, such approaches often rely on externally provided step-by-step guidance, limiting their scalability in open-world environments where agents are expected to interpret high-level goals and navigate autonomously Hong et al. (2025). Object Goal Navigation (ObjectNav) presents a compelling alternative by requiring agents to locate target objects based on semantic cues, without depending on dense instruction sequences. While ObjectNav has been widely studied in ground-based indoor scenarios Kolve et al. (2017); Deitke et al. (2020); Ramakrishnan et al. (2021), its potential in aerial navigation across large-scale, unstructured outdoor environments remains underexplored. Bridging this gap could open new avenues for deploying autonomous aerial systems capable of goal-driven exploration in complex real-world settings.
@@ 2.1 Object Goal Navigation
ObjectNav tasks Anderson et al. (2018) require an agent to navigate towards a specified object category based solely on egocentric visual observations. Early benchmarks such as AI2-THOR  Kolve et al. (2017) and Gibson  Xia et al. (2018) provided the initial testbeds for studying embodied navigation in indoor environments. Later, RoboTHOR  Deitke et al. (2020) and Habitat-Matterport 3D Dataset (HM3D)  Ramakrishnan et al. (2021) introduced more scalable and realistic simulation platforms, establishing standardized evaluation protocols for ObjectNav tasks. Methodologically, early works employed end-to-end reinforcement learning or imitation learning  Ye and Yang (2021); Ye et al. (2021); Ramrakhya et al. (2023), followed by approaches incorporating semantic mapping and modular planning pipelines  Chaplot et al. (2020); Luo et al. (2022); Campari et al. (2022), and more recently, zero-shot ObjectNav leveraging Vision-Language Models (VLMs) to generalize towards unseen object categories  Zhao et al. (2023b, a); Chen et al. (2023). However, most prior efforts have focused on small-scale, static, and semantically homogeneous indoor environments with predominantly planar navigation. In contrast, we extend ObjectNav to large-scale, outdoor UAV scenarios, where agents operate in open-world landscapes, such as urban areas, forests, and water bodies, with three-dimensional movement and diverse semantic targets, presenting fundamentally new challenges in perception, planning, and generalization.
@@ 2.2 Aerial Navigation
Aerial navigation has attracted increasing interest in recent years, with early efforts centered on conventional tasks such as map-based planning and visual navigation Lu et al. (2018); Ait Saadi et al. (2022); Jones et al. (2023). More recently, the field has moved toward VLN, aiming to enable UAVs to understand and act on natural language instructions in large-scale, open-world environments. Notable contributions include AerialVLN Liu et al. (2023a), the first UAV-centric VLN benchmark, as well as subsequent works like GeoText-1652 Chu et al. (2024a) and CityNav Lee et al. (2024), which focus on fine-grained geospatial grounding and the integration of geographic context. Simulation platforms such as TravelUAV Wang et al. (2024b) and OpenFly Gao et al. (2025) have further advanced the field by providing scalable environments and extensive trajectory datasets, while methods like STMR Gao et al. (2024), GridVLN Zhao et al. (2025a), NavAgent Liu et al. (2024), and GeoNav Xu et al. (2025) have introduced sophisticated representations for spatial reasoning. Existing aerial VLN approaches often rely on discrete waypoint execution or teleport-based movement, thus bypassing critical aspects of real-world navigation, such as obstacle avoidance, motion continuity, and collision handling Yue et al. (2024). In contrast, we formulate a task where UAVs execute parameterized, physically grounded actions to explore open-world environments and locate semantic targets based on high-level instructions. This setup removes dependence on step-by-step guidance and better captures the physical and semantic complexities of real-world UAV operations, supporting more realistic evaluation of autonomous aerial systems.
@@
GeoText Chu et al. (2024a) Aerial VLN Location Movement Instruction
AerialVLN Liu et al. (2023a) Aerial VLN Location Movement Instruction
@@ 3 Task Definition
UAV-ON defines an instance-level ObjectNav task in open-world settings, where UAVs navigate based on semantic instructions. At the beginning of each episode, the UAV is randomly initialized with a 6-DoF pose $P_{0}=[x,y,z,0,\psi,0]$, where $(x,y,z)\in\mathbb{R}^{3}$ denotes its position and $\psi$ is the yaw angle. The agent receives a semantic instruction $c=\{\text{name, size, description}\}\in\mathcal{C}$, which includes the target object’s category, an estimated size, and an instance-level visual description. The goal object is guaranteed to lie within a horizontal search radius $r=50$ units centered at the starting location. With only multi-view RGB-D cameras onboard, the UAV performs navigation without GPS or global maps, relying entirely on egocentric perception. The agent is expected to understand the semantic instruction, infer object-scene relevance, and explore efficiently using only onboard sensory inputs. The action space consists of continuous, parameterized controls for translation, rotation, and stopping. Each episode terminates when the agent outputs a Stop command, collides with an obstacle, or reaches the maximum step limit of 150. Following Liu et al. (2023a), a success episode is defined by the agent stopping within 20 units of the target object.
@@ 4.2 Action Space
Unlike prior aerial navigation benchmarks that rely on fixed-step, discrete action spaces, this formulation provides a continuous control interface that more accurately reflects the real-world dynamics of UAVs. Actions are physically executed rather than teleported, requiring agents to ensure each selected motion yields a collision-free trajectory. Any contact with obstacles is considered a failure, significantly raising the standards for safe and reliable navigation. The setup supports fixed-step motion strategies Liu et al. (2023a) while also enabling adaptive, context-sensitive behaviors in challenging environments with clutter or semantic ambiguity. With fine-grained motion control, the agent can more precisely adapt to variations in object proximity, obstacle density, and goal ambiguity. Its pose is continuously adjusted according to the semantics of the chosen action and its control magnitude, enabling seamless integration between perception and action.
@@ 5.2 Dataset Analysis
UAV-ON includes 14 high-fidelity outdoor environments with diverse geographic structures, ranging from compact urban parks to expansive mountainous terrains. Each scene contains multiple semantically labeled regions and varies in horizontal scale (e.g., from 350×250350$\times$ 250350 × 250 to 1400×12501400$\times$ 12501400 × 1250 units), resulting in rich spatial and structural diversity for aerial navigation. Across all scenes, we place a total of 1270 unique target objects over a total horizontal area of approximately 9 million square units, yielding an average object density of about 1.4 per 100 square units. Each object is paired with a semantic instruction that enables agents to reason about both appearance and contextual relevance. To illustrate the variety of target categories and linguistic expressions, we generate word clouds from the name and description fields (see Figure 2), which highlight the breadth of concepts and cues used to guide navigation. The integration of realistic, large-scale environments with semantically grounded targets and linguistically rich instructions makes UAV-ON a challenging and comprehensive benchmark for aerial ObjectNav.
@@ 5.3 Dataset Split
We construct the training set using 10 diverse outdoor environments, generating a total of 10,000 navigation episodes. The number of tasks per environment is proportionally allocated based on spatial size and object distribution to ensure balanced coverage across different scene types. Each episode is initialized at a random location and paired with a reachable target within a fixed search radius, forming a local navigation context. During training, the UAV can access its absolute position and the real-time Euclidean distance to the target, allowing reinforcement learning  Sutton et al. (1998) methods to optimize navigation strategies based on egocentric observations and distance-based rewards. Furthermore, we voxelize each scene into 1-unit 3D grids and apply a three-dimensional A* algorithm  Hart et al. (1968) to compute the shortest traversable path from the agent to the target. These paths provide safe, expert trajectories that can be used to support imitation learning approaches  Schaal (1996). Our training environments form a semantically diverse and spatially rich search space, offering strong support for both interaction-based and demonstration-based learning methods.
@@ 5.4 Metrics
Here, $l_{i}$ is the shortest geodesic distance between the start and the goal, computed by discretizing the environment into a 1-unit occupancy grid and applying the A* Hart et al. (1968) algorithm. $p_{i}$ denotes the actual trajectory length of the agent in episode $i$.
@@ 6.1 Baselines
Aerial ObjectNav Agent (AOA): We introduce a zero-shot navigation system that taps into the spatial reasoning and exploration capabilities of LLMs within complex, open-world 3D settings. At each timestep, the agent receives RGB and depth inputs from four fixed perspectives — front, left, right, and downward — paired with a structured target prompt that describes the object’s name, estimated size, and visual characteristics. RGB frames are transformed into textual descriptions using a pre-trained Qwen-VL model Wang et al. (2024a), while depth data is distilled into compact $\times 3$ 3 spatial matrices through min pooling. The system also logs the last 10 poses as 4D vectors $[x,y,z,\psi]$, offering contextual cues on recent movements.
All these multimodal inputs—including target semantics, image-language descriptions, geometric depth maps, pose history, and task constraints—are integrated into a structured prompt, which is subsequently passed to GPT-4o mini Achiam et al. (2023). Functioning as a high-level controller, GPT generates a semantic action command along with a continuous control magnitude, such as “move forward 3 units” or “rotate left 60 degrees.” To facilitate comparative evaluation, we develop two variants of the AOA framework. AOA (Fixed-step) employs constant translation and rotation parameters, mirroring previous UAV navigation baselines Liu et al. (2023a); Lee et al. (2024), thus ensuring compatibility with discrete control benchmarks. By contrast, AOA (Variable-step) exploits the reasoning ability of GPT to flexibly determine both the action type and its corresponding magnitude, aligning more naturally with the benchmark’s continuous action space and allowing for greater behavioral adaptability. Both AOA variants operate entirely in a zero-shot regime without any model training or fine-tuning, offering a standardized framework for evaluating the spatial generalization capabilities of LLMs.To ensure navigational validity, we implement a post-processing mechanism that replaces out-of-bound movement actions with in-place rotations, allowing the UAV to reorient itself without exiting the scene. This adjustment was motivated by early observations that GPT occasionally exhibits limited spatial awareness. We do not enforce a minimum flight altitude; the system relies on onboard depth cameras to avoid low-altitude collisions.
@@ Appendix B Instance-level Object Prompt Construction
To support precise and semantically grounded object-goal navigation, we construct structured instance-level prompts that provide detailed natural language descriptions of target objects. These prompts are used to communicate high-level task goals to the navigation agent in a format compatible with large language models. As illustrated in Figure 5, the prompt construction process begins with multi-view observations of each object instance, captured within the simulation environment. Specifically, we collect front, plan, and elevation views, which together provide comprehensive visual and spatial context. This multi-view input helps capture key features such as silhouette, dimensions, and placement relative to the ground plane—elements critical for downstream recognition and localization. We then leverage GPT-4 to generate descriptive prompts based on these visual observations. Each prompt consists of three components: (1) a visual category label (e.g., Statue, Chair), (2) an estimated size in plan view (e.g., mid(2.75$\times$2.5=6.79 unit squares)), and (3) a fine-grained natural language description that captures the object’s shape, material, and semantic attributes (e.g., Organic seated figure, oxidized bronze and dark stone, commemorative sculptural monument). These detailed instance-level prompts are then provided to the navigation model as the semantic goal input. This approach ensures that the agent receives a rich, human-interpretable description that supports goal understanding, spatial reasoning, and object disambiguation. To facilitate reproducibility and scalability, we release a lightweight, open-source tool for automated prompt generation. Given input images, the tool formats them appropriately and interfaces with GPT to produce structured object descriptions. The toolkit and usage instructions are publicly available at: https://chatgpt.com/g/g-6819b2befeac81918550152db018a5dd-object-description-generator.
@@ Appendix C Justification for the Maximum Exploration Step Limit
To define a principled upper bound on the number of exploration steps in the UAV-ON task, we analyze the perceptual geometry of the UAV and the spatial extent of the search region. Each UAV is equipped with four RGB-D cameras, each with a field of view (FOV) of $90^{\circ}\times 90^{\circ}$. For typical small to medium-sized objects, we assume that the downward-facing camera provides sufficient coverage when the UAV hovers at an altitude of 5 meters, enabling reliable ground-level visibility without excessive occlusion (see Figure 6(a)). Given this sensor configuration, we construct a discrete 2D grid map over the search region, where each cell approximates the area observable in a single downward view from a distance of 5 meters. The search region is defined as a circular area with a radius of 50 units in the $X-Y$ plane, centered at the UAV’s starting location. As visualized in Figure 6(b), all grid cells whose centers or partial areas fall within this circle are marked as part of the valid search space. This results in a total of 88 grid cells to be visited for complete coverage of the region. To estimate a lower bound on the number of steps required for full exploration, we simulate a deterministic zig-zag traversal strategy using only discrete translational actions: move forward, move left, and move right. This policy does not utilize global planning or semantic priors, but rather reflects a basic spatial sweep approach. As shown in Figure 6(c), visiting all 88 valid grid cells requires 92 steps under idealized conditions. To account for practical factors such as the initial movement from the center, minor altitude adjustments, and occasional maneuvers for obstacle avoidance, we conservatively increase this estimate by a factor of 1.5. Based on these considerations, we set the maximum exploration step limit to 150. This allows sufficient action budget for agents to conduct exhaustive local exploration, while also accommodating more reactive or adaptive behaviors. Importantly, this upper bound ensures a fair evaluation setting across methods by providing enough flexibility to complete the task without artificially limiting navigation capacity.

### v2

@@ 1 Introduction
Recent progress in aerial Vision-and-Language Navigation (VLN) Liu et al. (2023); Chu et al. (2024); Wang et al. (2024b); Gao et al. (2025); Zhao et al. (2025) has been largely driven by detailed, fine-grained language instructions. However, such approaches often rely on externally provided step-by-step guidance, limiting their scalability in open-world environments where agents are expected to interpret high-level goals and navigate autonomously Hong et al. (2025). Object Goal Navigation (ObjectNav) presents a compelling alternative by requiring agents to locate target objects based on semantic cues, without depending on dense instruction sequences. While ObjectNav has been widely studied in ground-based indoor scenarios Kolve et al. (2017); Deitke et al. (2020); Ramakrishnan et al. (2021), its potential in aerial navigation across large-scale, unstructured outdoor environments remains underexplored. Bridging this gap could open new avenues for deploying autonomous aerial systems capable of goal-driven exploration in complex real-world settings.
@@ 2.1 Object Goal Navigation
ObjectNav tasks Anderson et al. (2018) require an agent to navigate towards a specified object category based solely on egocentric visual observations. Early benchmarks such as AI2-THOR  Kolve et al. (2017) and Gibson  Xia et al. (2018) provided the initial testbeds for studying embodied navigation in indoor environments. Later, RoboTHOR  Deitke et al. (2020) and Habitat-Matterport 3D Dataset (HM3D)  Ramakrishnan et al. (2021) introduced more scalable and realistic simulation platforms, establishing standardized evaluation protocols for ObjectNav tasks. Methodologically, early works employed end-to-end reinforcement learning or imitation learning  Ye and Yang (2021); Ye et al. (2021); Ramrakhya et al. (2023), followed by approaches incorporating semantic mapping and modular planning pipelines  Chaplot et al. (2020); Luo et al. (2022); Campari et al. (2022), and more recently, zero-shot ObjectNav leveraging Vision-Language Models (VLMs) to generalize towards unseen object categories  Zhao et al. (2022, 2023); Chen et al. (2023). However, most prior efforts have focused on small-scale, static, and semantically homogeneous indoor environments with predominantly planar navigation. In contrast, we extend ObjectNav to large-scale, outdoor UAV scenarios, where agents operate in open-world landscapes, such as urban areas, forests, and water bodies, with three-dimensional movement and diverse semantic targets, presenting fundamentally new challenges in perception, planning, and generalization.
@@ 2.2 Aerial Navigation
Aerial navigation has attracted increasing interest in recent years, with early efforts centered on conventional tasks such as map-based planning and visual navigation Lu et al. (2018); Ait Saadi et al. (2022); Jones et al. (2023). More recently, the field has moved toward VLN, aiming to enable UAVs to understand and act on natural language instructions in large-scale, open-world environments. Notable contributions include AerialVLN Liu et al. (2023), the first UAV-centric VLN benchmark, as well as subsequent works like GeoText-1652 Chu et al. (2024) and CityNav Lee et al. (2024), which focus on fine-grained geospatial grounding and the integration of geographic context. Simulation platforms such as TravelUAV Wang et al. (2024b) and OpenFly Gao et al. (2025) have further advanced the field by providing scalable environments and extensive trajectory datasets, while methods like STMR Gao et al. (2024), GridVLN Zhao et al. (2025), NavAgent Liu et al. (2024), and GeoNav Xu et al. (2025) have introduced sophisticated representations for spatial reasoning. Existing aerial VLN approaches often rely on discrete waypoint execution or teleport-based movement, thus bypassing critical aspects of real-world navigation, such as obstacle avoidance, motion continuity, and collision handling Yue et al. (2024). In contrast, we formulate a task where UAVs execute parameterized, physically grounded actions to explore open-world environments and locate semantic targets based on high-level instructions. This setup removes dependence on step-by-step guidance and better captures the physical and semantic complexities of real-world UAV operations, supporting more realistic evaluation of autonomous aerial systems.
@@
GeoText Chu et al. (2024) Aerial VLN Location Movement Instruction
AerialVLN Liu et al. (2023) Aerial VLN Location Movement Instruction
@@ 3 Task Definition
UAV-ON defines an instance-level ObjectNav task in open-world settings, where UAVs navigate based on semantic instructions. At the beginning of each episode, the UAV is randomly initialized with a 6-DoF pose $P_{0}=[x,y,z,0,\psi,0]$, where $(x,y,z)\in\mathbb{R}^{3}$ denotes its position and $\psi$ is the yaw angle. The agent receives a semantic instruction $c=\{\text{name, size, description}\}\in\mathcal{C}$, which includes the target object’s category, an estimated size, and an instance-level visual description. The goal object is guaranteed to lie within a horizontal search radius $r=50$ units centered at the starting location. With only multi-view RGB-D cameras onboard, the UAV performs navigation without GPS or global maps, relying entirely on egocentric perception. The agent is expected to understand the semantic instruction, infer object-scene relevance, and explore efficiently using only onboard sensory inputs. The action space consists of continuous, parameterized controls for translation, rotation, and stopping. Each episode terminates when the agent outputs a Stop command, collides with an obstacle, or reaches the maximum step limit of 150. Following Liu et al. (2023), a success episode is defined by the agent stopping within 20 units of the target object.
@@ 4.2 Action Space
Unlike prior aerial navigation benchmarks that rely on fixed-step, discrete action spaces, this formulation provides a continuous control interface that more accurately reflects the real-world dynamics of UAVs. Actions are physically executed rather than teleported, requiring agents to ensure each selected motion yields a collision-free trajectory. Any contact with obstacles is considered a failure, significantly raising the standards for safe and reliable navigation. The setup supports fixed-step motion strategies Liu et al. (2023) while also enabling adaptive, context-sensitive behaviors in challenging environments with clutter or semantic ambiguity. With fine-grained motion control, the agent can more precisely adapt to variations in object proximity, obstacle density, and goal ambiguity. Its pose is continuously adjusted according to the semantics of the chosen action and its control magnitude, enabling seamless integration between perception and action.
@@ 5.2 Dataset Analysis
UAV-ON includes 14 high-fidelity outdoor environments with diverse geographic structures, ranging from compact urban parks to expansive mountainous terrains. Each scene contains multiple semantically labeled regions and varies in horizontal scale (e.g., from $350\times 250$ to $1400\times 1250$ units), resulting in rich spatial and structural diversity for aerial navigation. Across all scenes, we place a total of 1270 unique target objects over a total horizontal area of approximately 9 million square units, yielding an average object density of about 1.4 per 100 square units. Each object is paired with a semantic instruction that enables agents to reason about both appearance and contextual relevance. To illustrate the variety of target categories and linguistic expressions, we generate word clouds from the name and description fields (see Figure 2), which highlight the breadth of concepts and cues used to guide navigation. The integration of realistic, large-scale environments with semantically grounded targets and linguistically rich instructions makes UAV-ON a challenging and comprehensive benchmark for aerial ObjectNav.
@@ 5.3 Dataset Split
We construct the training set using 10 diverse outdoor environments, generating a total of 10,000 navigation episodes. The number of tasks per environment is proportionally allocated based on spatial size and object distribution to ensure balanced coverage across different scene types. Each episode is initialized at a random location and paired with a reachable target within a fixed search radius, forming a local navigation context. During training, the UAV can access its absolute position and the real-time Euclidean distance to the target, allowing reinforcement learning  Sutton and Barto (1998) methods to optimize navigation strategies based on egocentric observations and distance-based rewards. Furthermore, we voxelize each scene into 1-unit 3D grids and apply a three-dimensional A* algorithm  Hart et al. (1968) to compute the shortest traversable path from the agent to the target. These paths provide safe, expert trajectories that can be used to support imitation learning approaches  Schaal (1996). Our training environments form a semantically diverse and spatially rich search space, offering strong support for both interaction-based and demonstration-based learning methods.
@@ 5.4 Metrics
Here, $l_{i}$ is the shortest geodesic distance between the start and the goal, computed by discretizing the environment into a 1-unit occupancy grid and applying the A* Hart et al. (1968) algorithm. $p_{i}$ denotes the actual trajectory length of the agent in episode ii.
@@ 6.1 Baselines
Aerial ObjectNav Agent (AOA): We introduce a zero-shot navigation system that taps into the spatial reasoning and exploration capabilities of LLMs within complex, open-world 3D settings. At each timestep, the agent receives RGB and depth inputs from four fixed perspectives — front, left, right, and downward — paired with a structured target prompt that describes the object’s name, estimated size, and visual characteristics. RGB frames are transformed into textual descriptions using a pre-trained Qwen-VL model Wang et al. (2024a), while depth data is distilled into compact $3\times 3$ spatial matrices through min pooling. The system also logs the last 10 poses as 4D vectors $x,y,z,\psi]$, offering contextual cues on recent movements.
All these multimodal inputs—including target semantics, image-language descriptions, geometric depth maps, pose history, and task constraints—are integrated into a structured prompt, which is subsequently passed to GPT-4o mini OpenAI et al. (2023). Functioning as a high-level controller, GPT generates a semantic action command along with a continuous control magnitude, such as “move forward 3 units” or “rotate left 60 degrees.” To facilitate comparative evaluation, we develop two variants of the AOA framework. AOA (Fixed-step) employs constant translation and rotation parameters, mirroring previous UAV navigation baselines Liu et al. (2023); Lee et al. (2024), thus ensuring compatibility with discrete control benchmarks. By contrast, AOA (Variable-step) exploits the reasoning ability of GPT to flexibly determine both the action type and its corresponding magnitude, aligning more naturally with the benchmark’s continuous action space and allowing for greater behavioral adaptability. Both AOA variants operate entirely in a zero-shot regime without any model training or fine-tuning, offering a standardized framework for evaluating the spatial generalization capabilities of LLMs.To ensure navigational validity, we implement a post-processing mechanism that replaces out-of-bound movement actions with in-place rotations, allowing the UAV to reorient itself without exiting the scene. This adjustment was motivated by early observations that GPT occasionally exhibits limited spatial awareness. We do not enforce a minimum flight altitude; the system relies on onboard depth cameras to avoid low-altitude collisions.
@@ Appendix B Instance-level Object Prompt Construction
[…]
@@ Appendix C Justification for the Maximum Exploration Step Limit
To define a principled upper bound on the number of exploration steps in the UAV-ON task, we analyze the perceptual geometry of the UAV and the spatial extent of the search region. Each UAV is equipped with four RGB-D cameras, each with a field of view (FOV) of $90^{\circ}\times 90^{\circ}$. For typical small to medium-sized objects, we assume that the downward-facing camera provides sufficient coverage when the UAV hovers at an altitude of 5 meters, enabling reliable ground-level visibility without excessive occlusion (see Figure 6(a)). Given this sensor configuration, we construct a discrete 2D grid map over the search region, where each cell approximates the area observable in a single downward view from a distance of 5 meters. The search region is defined as a circular area with a radius of 50 units in the XYX-Y plane, centered at the UAV’s starting location. As visualized in Figure 6(b), all grid cells whose centers or partial areas fall within this circle are marked as part of the valid search space. This results in a total of 88 grid cells to be visited for complete coverage of the region. To estimate a lower bound on the number of steps required for full exploration, we simulate a deterministic zig-zag traversal strategy using only discrete translational actions: move forward, move left, and move right. This policy does not utilize global planning or semantic priors, but rather reflects a basic spatial sweep approach. As shown in Figure 6(c), visiting all 88 valid grid cells requires 92 steps under idealized conditions. To account for practical factors such as the initial movement from the center, minor altitude adjustments, and occasional maneuvers for obstacle avoidance, we conservatively increase this estimate by a factor of 1.5. Based on these considerations, we set the maximum exploration step limit to 150. This allows sufficient action budget for agents to conduct exhaustive local exploration, while also accommodating more reactive or adaptive behaviors. Importantly, this upper bound ensures a fair evaluation setting across methods by providing enough flexibility to complete the task without artificially limiting navigation capacity.